How many cubic centimeters of hydrogen chloride gas at s.t.p would be required to precipitate all the silver ions from 32cm 0.08 mole silver nitrate solution
1
Expert's answer
2017-04-28T10:16:08-0400
AgNO3 + HCl = AgCl + HNO3
CM = n/V ; n = CM · V
According to equestion: n (AgNO3) = n (HCl)
0.032 · 0.08 = 0.08 · V (HCl)
V (HCl) = (0.032 · 0.08)/0.08 = 0.032 mol